### v2

# Surface-Aware Distilled 3D Semantic Features

Abstract.
Many 3D tasks such as pose alignment, animation, motion transfer, and 3D reconstruction rely on establishing correspondences between 3D shapes. This challenge has recently been approached by pairwise matching of semantic features from pre-trained vision models. However, despite their power, these features struggle to differentiate instances of the same semantic class such as “left hand” versus “right hand” which leads to substantial mapping errors. To solve this, we learn a surface-aware embedding space that is robust to these ambiguities while facilitating shared mapping for an entire family of 3D shapes. Importantly, our approach is self-supervised and requires only a small number of unpaired training meshes to infer features for new possibly imperfect 3D shapes at test time. We achieve this by introducing a contrastive loss that preserves the semantic content of the features distilled from foundational models while disambiguating features located far apart on the shape’s surface. We observe superior performance in correspondence matching benchmarks and enable downstream applications including 2D-to-3D and 3D-to-3D texture transfer, in-part segmentation, pose alignment, and motion transfer in low-data regimes. Unlike previous pairwise approaches, our solution constructs a joint embedding space, where both seen and unseen 3D shapes are implicitly aligned without further optimization. The code is available at https://graphics.tudelft.nl/SurfaceAware3DFeatures.
Semantic Features, Contrastive Learning, Motion Transfer, Reposing, Shape Correspondences
$c$s: Computing methodologies Shape analysisccs: Computing methodologies Dimensionality reduction and manifold learningccs: Computing methodologies Motion processing
Figure 1. We introduce a surface-aware feature embedding space based on 2D pre-trained foundational models. In contrast to related works, our embedding space separates instances of the same semantic class (e.g. right vs. left instances for “hand/paw”) which facilitates many downstream applications: a) Texturing of even incomplete 3D shapes based on a 2D image (the input image was produced by ChatGPT), b) 3D correspondences between non-isometric shapes, c) Re-posing of meshes based on a single source and target pair, d) Pose alignment of 3D meshes with dense and sparse point correspondences. :
1. Introduction
Mapping 3D shapes into a shared space guaranteeing mutual correspondences is important for many applications, including 3D registration, pose alignment, motion transfer, as well as static and dynamic 3D reconstruction. Historically, geometric descriptors have been used to determine matches between pairs of 3D shapes under isometric deformations, but they struggle with non-isometric deformations (Sun et al., 2009; Tombari et al., 2010; Aubry et al., 2011). In contrast, neural features, stemming from pre-trained 2D vision models, have recently achieved great success in identifying correspondences between pairs of vastly different shapes (Tang et al., 2023; Luo et al., 2023; Zhang et al., 2024b; Wimmer et al., 2024), such as mapping from cats to lions. In this paper, we make another important step by moving from one-to-one pairwise shape correspondence matching to a joint embedding space establishing many-to-many shape correspondences.
Despite their inter-class robustness, neural features often struggle to disambiguate between instances of the same class like “left hand” and “right hand” (see  Fig. 3). Such mismatches can lead to substantial errors in downstream applications (see Sec. 6). Recent research has demonstrated that these features contain global pose information and that disambiguation is possible in a 2D scenario (Zhang et al., 2024a). However, achieving the same effect on distilled 3D features is not trivial, especially in a low-data regime, which is prevalent in 3D, where data acquisition and labeling is difficult.
Our work improves 3D neural features distilled from pre-trained 2D vision models by embedding them into a space disambiguating intraclass instances. We achieve this without large annotated datasets using a self-supervised learning scheme guided by in-shape geodesic distances without the need for shape pairs. Training with a limited number of 3D meshes, our method produces space of surface-aware features establishing multi-faceted correspondences for diverse new shapes without any further fine-tuning. In quantitative and qualitative comparisons to prior work, we demonstrate superior suitability of these features to serve as robust descriptors for matching and as building blocks for solving other tasks. Since geodesic distances are not used during inference, our method has only a minimal overhead from its shallow neural encoder and its point-wise nature makes it robust to varying mesh complexity or shape incompleteness. Finally, the encoder preserves compatibility with per-pixel image features, and hence, also naturally establishes robust 2D-to-3D mappings.
In summary, we make the following contributions: 1. We introduce a novel contrastive loss for self-supervised distillation of 3D features. 2. We quantitatively demonstrate the effectiveness of our surface-aware features in pose transfer, correspondence matching and skinning weight regression. 3. We showcase versatility of our approach in additional downstream applications including pose alignment, instance-based part segmentation and 2D-to-3D or 3D-to-3D texture transfer. 4. We show the versatility of our features when matching many-to-many shapes of not only humanoids and animals but also other classes.
2. Related Works
Our method utilizes contrastive learning to embed semantic features from foundational models to a space enabling robust n-to-n 3D shape matching. In this section, we discuss prior work in these three areas.
2.1. Image-based features for 3D shapes
Image-based features emerge in large visual models for 2D image tasks. Self-supervised features from Vision Transformers, such as DINO-ViT (Caron et al., 2021) and DINOv2 (Oquab et al., 2024), locally encode semantic information useful for segmentation (Caron et al., 2021) or image-to-image correspondence matching (Amir et al., 2021). SD-DINO (Zhang et al., 2023a) adds complementary features from the diffusion-based image synthesis model Stable Diffusion (Rombach et al., 2022). Lifting these features to 3D has enabled the self-supervised construction of canonical surface maps (Shtedritski et al., 2024), transfer of appearance between 3D shapes (Fischer et al., 2024), 3D animation (Uzolas et al., 2024), keypoint detection (Wimmer et al., 2024) or matching of surface correspondences (Dutt et al., 2024; Morreale et al., 2024; Chen et al., 2025). However, despite their semantic versatility, disambiguating between intraclass instances, such as left and right hands, remains challenging but possible, as shown in a recent 2D image study (Zhang et al., 2024a). This motivates our 3D shape descriptors for resolving in- stance ambiguity.
Prior work tackled this ambiguity by mapping shapes to a spherical template (Mariotti et al., 2024), which is difficult for complex shapes including humans. Alternatively, Liu et al. (2025) recovered non-isometric correspondences from a 2D semantic flow learned from vision features. Instead, we adapt Diff3F features (Dutt et al., 2024) in 3D space and resolve the ambiguity through contrastive learning enforcing geodesic distances.
Geodesic distances have previously supported point cloud analysis (He et al., 2019) and more recently, NIE (Jiang et al., 2023) and the concurrent work DV-Matcher (Chen et al., 2025) similarly utilize geodesic distances for feature disambiguation. Yet, the previously mentioned methods rely on aligned mesh pairs or a learned alignment, while our method learns purely from intrinsic properties of individual shapes. This eases adaptation to less common classes beyond humanoids and animals (Fig. 8) and cross-class mappings (Fig. 9).
Beyond vision-only models, multimodal large language models have recently been effective in image and 3D shape analysis including keypoint labeling (Gong et al., 2024) and shape co-segmentation (Abdelreheem et al., 2023). In our work, we focus on vision-only models because of their simplicity, but we consider a model combination a promising research direction.
2.2. Contrastive Learning
Contrastive learning embeds similar samples close to each other while keeping dissimilar samples apart. This can be achieved directly by minimizing and maximizing embedding distances for positive and negative pair samples, respectively (Chopra et al., 2005; Hadsell et al., 2006; Weinberger and Saul, 2009; Schroff et al., 2015) or indirectly, such as by optimizing performance in an auto-regressive task (Oord et al., 2018). Training pairs can be obtained by data augmentation (Chen et al., 2020), from memory banks (Wu et al., 2018; He et al., 2020), or by clustering (Caron et al., 2018, 2020). Learning with cross-domain labels yields joint embeddings, as demonstrated by CLIP (Radford et al., 2021) for text and images. Contrastive learning was applied to learn end-to-end pose transfer from multiple unregistered meshes of the same identity in different poses (Sun et al., 2023a). We design our contrastive loss to disambiguate intra-class instances guided by a geodesic metric, while learning from intrinsic properties of individual meshes rather than same-identity shape pairs.
2.3. Shape correspondences
Point-to-Point
Classical shape registration methods directly minimize global (Besl and McKay, 1992) or local (Brown and Rusinkiewicz, 2007) inter-shape distances making them susceptible to local minima (Yang et al., 2015). This motivates the design of more informative local geometric descriptors (Sun et al., 2009; Tombari et al., 2010; Aubry et al., 2011). These can alternatively be learned (Corman et al., 2014; Guo et al., 2015) from voxelized patches (Zeng et al., 2017; Gojcic et al., 2019; Attaiki et al., 2023) or from point clouds (Elbaz et al., 2017; Yew and Lee, 2018; Deng et al., 2018b, a, 2023). The learning can be supervised by labels (Corman et al., 2014) or achieved without them (Elbaz et al., 2017; Groueix et al., 2018; Lang et al., 2021; Zeng et al., 2021). Our method falls into the latter category, as our contrastive loss motivates our encoder to separate instances by approximating geodesic distances (Xia et al., 2021) without training data labels. This is conceptually similar to previous methods for near-isometric shape deformations (Mémoli and Sapiro, 2005; Halimi et al., 2019; Shamai and Kimmel, 2017). However, we distinctly do not measure geodesic distortions between shape pairs, and therefore we do not limit our method to isometric deformations, and we do not compute any geodesics during inference. Instead, we only use the geodesics to disambiguate information already available in the image-based features, which is critical for our results.
The correspondences can be recovered from descriptors by a matching (Fischler and Bolles, 1981), directly regressed (Lu et al., 2019; Wang and Solomon, 2019) or established on parametric templates (Groueix et al., 2018; Deprelle et al., 2019). Here, we focus on the descriptors themselves, and we show several different application scenarios in Sec. 6.
Surface mapping
Functional Maps (FMs) (Ovsjanikov et al., 2012) allow for matching on a surface. FMs are real-valued surface functions in the space of Laplace-Bertrami eigenfunctions, supporting linear transformations between shapes. Constrained to match surface descriptors for each shape (Sun et al., 2009; Tombari et al., 2010; Aubry et al., 2011) they allow extracting point-wise correspondences (Ovsjanikov et al., 2012; Rodolà et al., 2015). These functions can also be learned (Litany et al., 2017) often with little or no supervision (Roufosse et al., 2019; Donati et al., 2020; Ginzburg and Raviv, 2020; Halimi et al., 2019; Sun et al., 2023b). Extrinsic alignment can support nonisotropic deformations (Eisenberger et al., 2020a, b). In this work, we focus on improving features for direct point-to-point matching in the spatial domain, but we later demonstrate a combination of our features with FM.
Figure 2. Overview of our method. We feed images of a 3D shape rendered from multiple viewpoints to a pre-trained 2D vision model and extract features that are then projected back onto surface points $\mathbf{p}_{i}$ and aggregated into per-point features $\mathbf{f}_{i}$ (Sec. 3). Next, we pointwise embed the base features $\mathbf{f}_{i}$ into our surface-aware features $\mathbf{s}_{i}$ residing in a lower-dimensional space learned using our contrastive loss preserving geodesic distances $d_{i,j}$ and a reconstruction loss matching decoded features $\mathbf{\bar{f}}_{i}$ to $\mathbf{f}_{i}$ (Sec. 4). The surface-aware features $\mathbf{s}_{i}$ serve as robust descriptors for correspondence matching (Sec. 5) and base blocks for many down-stream applications (Sec. 6).
3. Preliminaries
We build upon methods that aggregate features from pre-trained 2D vision models on 3D meshes (Dutt et al., 2024; Wimmer et al., 2024; Morreale et al., 2024; Chen et al., 2025). In this section, we give a brief overview on these methods.
3.1. Reprojection of 2D Features
We represent a 3D shape as a triangular mesh with a tuple of $N$ vertices and $M$ triangular faces, that is, $\mathcal{M}{}:=(\{\mathbf{p}_{n}\in\mathbb{R}^{3}|n=1,...,N\},\{\mathbf{t}_{m}\in\mathbb{N}^{3}|m=1,...,M\})$. The rendering function $R_{rgb}:(\mathcal{M},\mathcal{C})\rightarrow\mathbf{I}_{rgb}$ projects $\mathcal{M}{}$ into a camera $\mathcal{C}$ and outputs an image $\mathbf{I}_{rgb}\in\mathbb{R}^{H\times W\times 3}$, with height $H$ and width $W$. Optionally, texturing is possible in $R_{rgb}(.)$ or as a ControlNet (Zhang et al., 2023b) post-processing. The image is then passed to a pre-trained vision model (Caron et al., 2021; Oquab et al., 2024; Rombach et al., 2022; Zhang et al., 2023a) to obtain dense semantic feature maps $\mathbf{F}\in\mathbb{R}^{h\times w\times f}$ with $h,w,f$ as two spatial and one feature dimension. Finally, per-vertex features $\mathbf{f}_{n}\in\mathbb{R}^{f}$ are obtained by projective texture mapping of $\mathbf{F}$ onto $\mathcal{M}$. To cover the whole surface, features are aggregated across multiple cameras, resulting in a set of features $\mathcal{F}_{\mathcal{M}}:=\{\mathbf{f}_{n}\in\mathbb{R}^{f}|n=1,...,N\}$. Throughout this work, we refer to $\mathcal{F}_{\mathcal{M}}$ as the base features on which our method is built. The exact choice of $\mathcal{F}_{\mathcal{M}}$ is orthogonal to our contribution but must encode semantic information. To this extent, we use Diff3F (Dutt et al., 2024) in this work.
Correspondence Matching
Features $\mathcal{F}_{\mathcal{M}}$ have been shown to encode strong semantic information useful for correspondence matching (Tang et al., 2023; Dutt et al., 2024). In the simplest case, the feature $\mathbf{f}_{n}\in\mathcal{F}_{\mathcal{T}}$ of a target mesh $\mathcal{T}$ that best matches the feature $\mathbf{f}_{m}\in\mathcal{F}_{\mathcal{S}}$ of a source mesh $\mathcal{S}$ is determined by maximizing the cosine similarity $\phi:\mathbb{R}^{f}\times\mathbb{R}^{f}\rightarrow\mathbb{R}$:
(1) $\phi(\mathbf{f}_{i},\mathbf{f}_{j})=\frac{\mathbf{f}_{i}^{T}\mathbf{f}_{j}}{\lVert\mathbf{f}_{i}\lVert_{2}\lVert\mathbf{f}_{j}\lVert_{2}},$
such that $\tau(\mathbf{p}_{m})=\arg\max_{\mathbf{p}_{n}}\phi(\mathbf{p}_{n}\to\mathbf{f}_{n},\mathbf{p}_{m}\to\mathbf{f}_{m})$ is the best matching point. However, the features $\mathcal{F}_{\mathcal{M}}$ do not differentiate between semantic instances well (see  Fig. 8) which we address by learning robust surface-aware features $\mathcal{S}_{\mathcal{M}}$.
4. Method
Our goal is to learn an embedding resolving instance ambiguities of the base features $\mathcal{F}_{\mathcal{M}}$ and obtain surface-aware features $\mathcal{S}_{\mathcal{M}}$ (see Fig. 2). We achieve this by training a point-based feature auto-encoder with a limited set of training meshes and our contrastive loss for self-supervision. At test time, we can produce surface-aware features for novel unseen shapes without additional fine-tuning.
4.1. Setup
Our method requires a potentially small set of training meshes $\mathbf{M}_{t}=\{\mathcal{M}{}_{i}|i=1,...,K\}$, each associated with base features $\mathcal{F}_{\mathcal{M}}$ obtained following Sec. 3 and normalized by a Euclidean norm such that $\forall\mathbf{f}_{n}\in\mathcal{F}_{\mathcal{M}},\lVert\mathbf{f}_{n}\lVert_{2}:=1$.
Unlike other approaches (Jiang et al., 2023; Deng et al., 2023), we do not require extrinsic canonical mesh alignment during training, because we rely only on intrinsic properties of individual meshes. Similar considerations were previously made for Functional maps (Ovsjanikov et al., 2012). Moreover, we inherit the rotation invariance of the base features demonstrated by Dutt et al. ([)Supplement “Robustness to Rotation”]dutt2024diffusion, although we observe a performance degradation if meshes are upside down and we avoid this in our inputs (see Appendix A.1).
Furthermore, our encoder is point-based and does not rely on shape completeness or a consistent topology. Both of these design choices favor generalization under transformations ranging from coordinate swap to shape reposing, and memory-unconstrained batch-based processing of even large shapes.
4.2. Separating Front Paw from Back Paw
Our embedding aims to separate multiple instances of the same class that are difficult to directly disambiguate in $\mathcal{F}_{\mathcal{M}}$. For example, consider the two surface points, $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$, on the bear’s paws in Fig. 2. The prevalent semantic significance of the “paw” concept hinders the separability of the corresponding base features $\mathbf{f}_{1}$ and $\mathbf{f}_{2}$. Fig. 3 illustrates this for human arms and Diff3F features (Dutt et al., 2024). To solve this, we train a point-wise feature autoencoder, producing our surface-aware features $\mathcal{S}_{\mathcal{M}}\subset\mathbb{R}^{s}$ in its embedding space. We motivate the feasibility of separation by the prior observations that vision features additionally carry information about the global pose (Zhang et al., 2024a). We postulate that this enables our model to distinguish between part instances when guided by their intrinsic distance. We adopt the geodesic distance $d_{1,2}$ between $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$ for this purpose. Following contrastive learning, we sample point pairs on a single shape to enforce $\phi(\mathbf{s}_{1},\mathbf{s}_{2})\approx d_{1,2}$ for $\mathbf{s}_{n}\in\mathcal{S}_{\mathcal{M}}$. We validate our choice of hyperspherical embedding space against Euclidean space in Sec. 5.5.
Model
We train a base feature encoder $\mathcal{E}(.)$, such that, following a normalization, $\mathbf{s}_{n}=\mathcal{E}(\mathbf{f}_{n})/\lVert\mathcal{E}(\mathbf{f}_{n})\lVert_{2}$ is a surface-aware feature $\mathbf{s}_{n}\in\mathbb{R}^{s}$ in a hypersphere embedding. During training, we randomly sample an unpaired training mesh $\mathcal{M}\in\mathbf{M}_{t}$ with base features $\mathcal{F}_{\mathcal{M}}$, which we encode pointwise to obtain $\mathcal{S}_{\mathcal{M}}$.
In each training iteration, we use furthest-point sampling to choose a random subset of $A$ anchor points $\mathbf{p}_{a}$ among the mesh vertices $\mathbf{p}_{i}\in\mathcal{M}{}$ and compute geodesic distances $d_{n,a}$ for each pair of a mesh and anchor point. We additionally rescale $d_{n,a}$ to a maximum of one, such that $d^{\prime}_{n,a}:=d_{n,a}/\max_{n,a}(d_{n,a})$, which removes the dependency on the scale of the mesh. We find this robust and leading to features later generalizing across morphologically equivalent shapes with different proportions (see an elephant vs. a giraffe in Fig. 8).
Subsequently, our contrastive loss preserves the rescaled geodesic distances in the embedding space:
(2) $\mathcal{L}_{c}=\frac{1}{NA}\sum_{n}^{N}\sum_{a}^{A}\left|d^{\prime}_{n,a}-\left(\frac{1-\phi(\mathbf{s}_{n},\mathbf{s}_{a})}{2}\right)\right|.$
This loss operates in a hyperspherical embedding space and utilizes cosine similarity mapped to the $[0,1]$ range. Hereby, $\mathcal{L}_{c}$ penalizes features close in the embedding space but distant on the shape surface and vice versa.
Furthermore, we found it beneficial for the preservation of semantic information to train a feature decoder $\mathbf{\bar{f}}_{n}=\mathcal{D}(\mathbf{s}_{n})/\lVert\mathcal{D}(\mathbf{s}_{n})\lVert_{2}$ in an autoencoder fashion. To this extent, we utilize a reconstruction loss:
(3) $\mathcal{L}_{r}=\frac{1}{N}\sum_{n}1-\phi(\mathbf{f}_{n},\mathbf{\bar{f}}_{n}).$
We train both the encoder and the decoder end-to-end with the combined loss $\mathcal{L}=w_{r}\mathcal{L}_{r}+w_{c}\mathcal{L}_{c}$, where a choice $w_{r}=w_{c}=1$ works well in our tests. We do not observe an increase in performance with a higher $w_{c}$.
Note that our training procedure, in contrast to related works (Lang et al., 2021; Deng et al., 2023; Chen et al., 2025), does not require target and source shape pairs.
4.3. Implementation
During preprocessing, we rasterize our triangular meshes and precompute base features for all vertices following Sec. 3. We implement our autoencoder in PyTorch2 (Ansel et al., 2024) and use the Polyscope renderer (Sharp et al., 2019b) for visualizations. The encoder $\mathcal{E}$ is a Multilayer Perceptron (MLP) consisting of three blocks, where each block has two linear layers, SiLU activation (Elfwing et al., 2018), and layer normalization (Ba et al., 2016). The first layer in each block employs a skip connection (He et al., 2016), while the second reduces the dimensionality by a factor of two. With Diff3F (Dutt et al., 2024) as base features, $\mathcal{E}$ reduces feature dimensionality from $f=2048$ to $s=256$. The decoder $\mathcal{D}$ is a mirrored copy of the encoder. We train our model on NVIDIA RTX 3090 for 50k iterations with the AdamW optimizer (Loshchilov and Hutter, 2017) and a learning rate of $0.0001$ which takes $\approx 2$ hours.
We choose an exponential moving average (Polyak and Juditsky, 1992) of the model with the lowest validation loss, without the need for any correspondence labels. Geodesic distances for training are calculated on the fly with the heat method (Crane et al., 2017) implemented in Geometry Central (Sharp et al., 2019a). No geodesics are required during inference and the computational cost is determined by the Diff3F baseline with a only a negligible overhead from our shallow encoder $\mathcal{E}$. For a shape with 10k vertices, this is less than 5 milliseconds on top of $\approx 4$ minutes from Diff3F. Moreover, downstream tasks cost benefits from the smaller feature dimensionality. Functional maps are calculated with the base algorithm (Ovsjanikov et al., 2012), provided by the Diff3F implementation (Dutt et al., 2024).
5. Experiments
Here, we first motivate the benefits of our surface-aware feature embedding space by visualizing its distribution. Next, we evaluate their effectiveness in tasks with quantitative benchmarks including pose transfer, skinning weight regression and 3D correspondence matching. Finally, we analyze the impact of our design choices in an ablation experiment.
Training
We train a single autoencoder on a joint dataset consisting of 49 animal samples from the SMAL dataset (Zuffi et al., 2017) and 49 humans from the SURREAL dataset (Groueix et al., 2018). We choose 2 samples from each dataset for validation. We use this single shared model without any additional optimization for all experiments, unless stated otherwise.
5.1. Exploration of Embedding Space
To illustrate the effect of our contrastive loss on feature separation, we compare the 2D projections of the Diff3F (Dutt et al., 2024) base features with our surface-aware features.
Figure 3. Two shapes (left) and a PCA-based 2D projections of their aggregated Diff3F base features and our surface-aware features (right). Notice the separation of limbs in our result compared to Diff3F. Our features originate from the same encoder for both shapes. The animal legs appear merged along the sagittal plane due limitations of the PCA projection, but they remain disambiguated in our feature space as demonstrated in Fig. 9.
Setup
We create two dataset SMPLeval and SMALeval unseen during training. The former consists of 50 randomly-sampled SMPL (Loper et al., 2023) shapes and poses from AMASS (Mahmood et al., 2019), while the latter consists of 50 randomly-sampled SMAL (Zuffi et al., 2017; Biggs et al., 2018) shapes in canonical poses. For each sample, we obtain the base features and surface-aware features as described in Sec. 3 and Sec. 4
Embedding
We project Diff3F features aggregated from SMPLeval to two dimensions using principal component analysis (PCA). In Fig. 3, we visualize the projection for two selected shapes from the same dataset. We repeat this with our surface-aware features. To avoid bias, we derive the visualized colors from the true SMPL (Loper et al., 2023) skinning weights $\mathbf{w}_{n}\in\mathbb{R}^{B}$ for both methods, where $B$ is the skinning weight dimension. We repeat this process for SMALeval. In Fig. 3, our method yields an interpretable embedding that separates the leg and hand instances for animals and humans despite not having access to extrinsic $(x,y,z)$ point positions. This validates the suitability of our features for downstream tasks and highlights the limitations of the Diff3F base features.
5.2. One-shot Pose Transfer
Figure 4. One-shot pose transfer using our features, Diff3F features, or Geometric descriptors. MSE $\times 10^{-4}$ is reported for human shapes.
We evaluate the performance of the surface-aware features in a one-shot re-posing task for arbitrary 2-manifold meshes. We use our features to fit a Neural Jacobian Field (NJF) (Aigerman et al., 2022) between poses of two input shapes and then apply it to re-pose a new target shape as described in Appendix B.
We sample 5 input shape pairs from SMPLeval by choosing one challenging pose as a target and one random shape as an initial pose. The remaining shape samples serve as test inputs for pose transfer with known ground truth. We report MSE for 240 such test pairs (see upper row, Fig. 4). Next, we repeat the same procedure with the base features and with the Geometric descriptors (Geo) (Aigerman et al., 2022), consisting of the face centroid, face normals, and a Wave-Kernel Signature (Aubry et al., 2011). Finally, we provide additional qualitative results for transfer of animal poses from TOSCA to SMALeval in Fig. 4.
We observe that our surface-aware features outperform the base features both quantitatively and qualitatively by correctly distinguishing individual posed limbs. The Geometric descriptors also perform well for humans, but they struggle with larger input and output shape differences in animals. Our features perform well in both cases.
[Uncaptioned image] SMPLeval SMALeval Ours .0014±.0006 .0048±.0008 Diff3F .0058±.0024 .0086±.0028 Geo .0210±.0031 .0287±.0017
Figure 5. Mean Squared Error of skinning weight regression ($\downarrow$ is better) and its distribution across the SMPL mesh surface.
5.3. Skinning Weight Regression
We train a simple regressor to predict skinning weights of a kinematic model, based on a single training sample.
Our pointwise regressor $\mathcal{W}_{\mathbf{s}}(\mathbf{s}_{n})$ consists of a linear layer and a Softmax activation and regresses skinning weights $\bar{\mathbf{w}}_{n}$ from our surface-aware features or, in case of $\mathcal{W}_{\mathbf{f}}(\mathbf{f}_{n})$ and $\mathcal{W}_{\mathbf{g}}(\mathbf{g}_{n})$, from the base features or the Geometric descriptors, respectively. We train all models fivefold supervised with the Mean Squared Error (MSE) and true weights separately on the SMPLeval and dsmalours datasets, and we report test MSE for the remaining unseen samples in the source datasets (see Fig. 5). Our features achieve lower errors and exhibit better robustness to instance ambiguities than the two alternatives. It is worth noting that the relatively lower dimensionality of the Geometric descriptors and base features affects the number of regressor parameters, impact of which is not studied in this experiment. However, our conclusions also hold for a 2-layer MLP regressor with an equal hidden dimensionality of 106 which matches the dimensionality of the Geometric descriptors.
5.4. Point-to-Point Correspondence Matching
Our features can be easily integrated into correspondence matching pipelines. Therefore, we replicate the evaluation setup of Diff3F (Dutt et al., 2024) and assess our surface-aware features in a correspondence matching task on human and animal shapes.
Data
For testing we use re-meshed versions of humans from SHREC’19 (Melzi et al., 2019; Donati et al., 2020) and animals from both SHREC’20 (Dyke et al., 2020) and the animal-only subset of TOSCA (Bronstein et al., 2008).
Baselines
We compare our method against the unsupervised image-based Diff3F method (Dutt et al., 2024), which also provides our base features, and against 3DCODED (Groueix et al., 2018), DPC (Lang et al., 2021) and SE-OrNet (Deng et al., 2023), which have been trained on thousands of samples, while our method is trained on less than 100 samples.
Metrics
We report commonly used point correspondence metrics for $1024$ points per mesh (Groueix et al., 2018; Lang et al., 2021; Deng et al., 2023; Dutt et al., 2024) [^1] We use the provided code and validate that we follow the same experimental procedure and metric definitions. . The correspondence error measures a distance between the computed correspondence point $\tau(\mathbf{p}_{n})$ (see Sec. 3) and the ground-truth correspondence point $\mathbf{t}_{n}^{gt}$ : $err=\frac{1}{n}\sum_{\mathbf{p}_{n}\in\mathcal{S}}\lVert\tau(\mathbf{p}_{n})-\mathbf{t}_{n}^{gt}\rVert_{2}^{2}$. The accuracy is the fraction of points with an error below a threshold $\epsilon\in[0,1]$: $acc(\epsilon)=\frac{1}{n}\sum_{\mathbf{p}_{n}\in\mathcal{S}}\mathbb{I}(\lVert\tau(\mathbf{p}_{n})-t_{n}^{gt}\rVert_{2}<\epsilon g)$, where $g$ is the maximal Euclidean distance in the target shape and $\mathbb{I}(.)$ is the indicator function.
Results
We provide quantitative results in Tbl. 1 and qualitative comparisons in Fig. 8. We find that our method achieves the lowest error on SHREC’19 and TOSCA, despite being trained on fewer samples than the supervised baselines. Furthermore, we outperform the Diff3F base features on both SHREC datasets in terms of accuracy. While Diff3F achieves a higher accuracy at 1% threshold in TOSCA, Fig. 8 shows that the accuracy of our model is higher for thresholds above $\approx 2$%. This suggests that our method excels in the removal of outliers that can be caused by mismatched components. Additionally, adapting our features to Functional maps (Ovsjanikov et al., 2012) (+ FM) distributes the error towards a lower mean at the cost of accuracy, and maintains the beneficial comparison to Diff3F+FM.
Fig. 8 shows that Diff3F struggles to separate intraclass instances such as left and right legs. In contrast, the results confirm the effectiveness of our contrastive loss in mitigating this issue. We observe the same behavior for SHREC’20 in Fig. 8 (top), which contains highly diverse animal shapes. Furthermore, our method generally produces visually smoother results (see Appendix D.1 and the supplementary videos on the project website).
Table 1. Comparison of our 3D correspondence matching to prior works 3DC (3D-CODED (Groueix et al., 2018)), DPC (Lang et al., 2021), SEN (SE-OrNet (Deng et al., 2023)), and Diff3F (Dutt et al., 2024). †) Numbers originate from (Dutt et al., 2024), *) Experiments were replicated, x) Omitted due to non-manifold meshes, + FM) Semantic features combined with Functional Maps. Accuracy is for the commonly used $\epsilon=1$ %. The per-column best results are bold and the second-to-best results are underlined.
SHREC’19 TOSCA SHREC’20
3DC† err ↓ 8.10 19.20 -
acc ↑ 2.10 0.50 -
DPC† err ↓ 6.26 3.74 2.13
acc ↑ 17.40 30.79 31.08
SEN† err ↓ 4.56 4.32 1.00
acc ↑ 21.41 33.25 31.70
Diff3F* err ↓ 1.69±1.44 4.51±5.48 5.34±10.22
acc ↑ 26.25±9.30 31.00±15.73 69.50±24.99
Diff3F + FM* err ↓ 1.51±1.65 x 4.44±7.87
acc ↑ 21.71±7.12 x 58.03±25.94
Ours err ↓ 0.43±0.76 1.65±2.15 3.89±8.90
acc ↑ 28.78±9.30 29.35±14.53 73.97±26.47
Ours + FM err ↓ 0.24±0.64 x 3.54±7.59
acc ↑ 24.83±6.80 x 63.61±24.34
Figure 6. Qualitative comparison on the SHREC’19 and TOSCA datasets with dense true correspondence labels provided by their authors. We show the source and target meshes with their ground truth correspondence labels (the two left-most columns) in comparison to correspondences computed using our surface-aware features (the forth column) and Diff3F base features (the right-most column). We further highlight the correspondence error on the mesh surface (the third and the fifth column). The error colormap is normalized per sample by the maximal error over both methods to keep the error scale comparable across columns but not across rows. Our surface-aware features notably improve separation of the limb instances.
Figure 7. Correspondence accuracy ($\uparrow$ is better) at different error distance thresholds for our own and Diff3F features.
Figure 8. Qualitative comparison of correspondence matching on TOSCA and ShapeNet  (Chang et al., 2015) (dense ground truth labels not available). Source shape (left) matched to target (right) using our and Diff3F features.
Other Shapes
Our method is applicable beyond humanoid and animal shapes, which we show by training two additional encoders for a subset of 50 chairs and 50 airplanes from ShapeNet (Chang et al., 2015). Here, we uniformly resample the mesh vertices for a better surface coverage (Wang et al., 2022).
For unseen shapes in Fig. 8, our surface-aware features again better distinguish same-class instances such as chair legs and airplane wings, supporting a wider applicability of our methodology. More examples are shown in Appendix D.2 and the supplementary videos.
5.5. Ablations
We motivate our design choices by ablation on various parts of our method in Tbl. 2 following the setup of Sec. 5.4.
Choice of Angular Space
We demonstrate the effectiveness of our hyperspherical embedding by replacing our contrastive loss $\mathcal{L}_{c}$ (Eq. 2) with three different options inspired by related work (see Appendix A.2). First, the Relative Geodesic Loss (RGL(Jiang et al., 2023) optimizes relative distances in a Euclidean embedding. Similarly, the Naive Geodesic Loss (NGL) minimizes absolute distances. Finally, the Geometrical Similarity Loss (GSL(Chen et al., 2025) enforces similarity of feature and surface distances in a local neighborhood. We remove feature and geodesic normalization wherever absolute magnitude needs to be learned. In Tbl. 2 (top), we observe that, except for correspondence accuracy for TOSCA, our contrastive loss $\mathcal{L}_{c}$ outperforms all of the alternatives in the correspondence matching task.
Contrastive and Reconstruction Loss
In Tbl. 2 (bottom), we individually assess our two losses. We see that the performance with only the reconstruction loss $\mathcal{L}_{r}$ is close to Diff3F. This indicates that the gain in performance does not originate predominantly from a smaller embedding space or from access to training data. Similarly, the contrastive loss $\mathcal{L}_{c}$ alone results in an accuracy drop compared to our full model. This justifies our autoencoder approach with both losses playing an import role. Ablations on the number of anchors can be found inAppendix A.2.
Table 2. Ablation on our method. Above the bar: Ablation on alternative losses inspired by related work (Chen et al., 2025; Jiang et al., 2023) compared to the unmodified Diff3F features. Below the bar: Our full method compared to its reduced variant omitting losses $\mathcal{L}_{c}$ or $\mathcal{L}_{r}$. The per-column best results are bold and the second-to-best results are underlined.
SHREC’19 TOSCA SHREC’20
RGL err ↓ 0.80±1.08 2.74±2.53 5.29±9.85
acc ↑ 20.16±10.03 16.53±10.41 55.23±21.93
NGL err ↓ 0.54±0.90 2.11±2.02 4.86±9.49
acc ↑ 18.84±9.59 18.86±11.50 58.97±23.12
GSL err ↓ 1.72±1.45 4.17±5.21 4.34±9.23
acc ↑ 26.89±9.07 29.77±14.92 73.39±26.31
Diff3F err ↓ 1.69±1.44 4.51±5.48 5.34±10.22
acc ↑ 26.25±9.30 31.00±15.73 69.50±24.99
only $\mathcal{L}_{r}$ err ↓ 1.65±1.44 4.70±5.64 4.87±9.38
acc ↑ 26.53±9.19 30.27±15.17 72.94±26.21
only $\mathcal{L}_{c}$ err ↓ 0.38±0.61 1.67±2.29 4.30±9.31
acc ↑ 26.21±8.78 25.58±13.88 70.08±25.17
Ours err ↓ 0.43±0.76 1.65±2.15 3.89±8.90
acc ↑ 28.78±9.30 29.35±14.53 73.97±26.47
6. Applications
We present additional downstream tasks that benefit from our surface-aware features learned in Sec. 5.
6.1. Instance-based Part Segmentation
Following the prior work (Dutt et al., 2024), we segment a target shape by clustering features around centroids from K-means clustering of source-shape features. In the top row of Fig. 9, we demonstrate a transfer from a big cat to a human and see that, unlike the Diff3F features, our surface-aware features disambiguate the limbs. In the bottom two rows, we repeat this experiment with a shared encoder trained on human, animals, and a subset of ShapeNet (see Appendix D.4) where a true mapping cannot be defined but our method finds reasonable analogies between the classes.
In Fig. 11, we repeat this with centroids obtained jointly from all SMPLeval and TOSCA samples. In contrast to Diff3F, our method successfully matches features across diverse shapes, which demonstrates our embedding’s capability of many-to-many shape matching without any additional pairwise optimization. Finally, we show similar results for chairs and airplanes in Fig. 12.
Figure 9. In the top row, 10 k-means cluster centers from the big cat were used to segment the human. In the bottom two rows, 8 k-means cluster centers from the animals were used to segment the chairs and airplanes with a shared encoder. Unlike Diff3F, our method successfully separates all limbs for a plausible mapping from animal limbs to human limbs, chair legs, or airplane wings.
Figure 10. Pose alignment of a source shape (color) into the pose of a target (gray). The boxes highlight challenging areas handled well by our method. For humans, we densely fit all the target vertices, while for animals, we only fit 55 % of the vertices as highlighted.
6.2. Pose Alignment
Our surface-aware features are also useful for pose alignment of a kinematic model to another 3D shape. To this end, we establish point correspondences between shape pairs as in Sec. 5.4 and optimize the kinematic pose parameters to minimize point-to-point distances (see Appendix C).
In Fig. 10, we align SMPLeval to SHREC’19, and SMALeval to DeformingThings4D (Li et al., 2021) animals. Benefiting from the robust instance separation, our method produces poses closer to the targets for both dense and sparse correspondences. See our video for a 3D shape animation obtained by aligning to a target shape sequence.
Figure 11. Results when clustering features across all samples in SMPLeval and TOSCA. Our method implicitly aligns semantically-related regions (shown as the same colors) across diverse 3D shapes in a self-supervised manner (the top row). Diff3F produces inconsistent labeling across different shape categories as well as lack of separability between instanced components such as individual limbs (the bottom row).
Figure 12. Results when clustering the surface-aware features across chairs (top row) and airplanes (bottom row) from ShapeNet (Chang et al., 2015). Note, that while we use a single shared encoder for all humanoid and animal shapes, we train a separate encoder for each ShapeNet class due to the large domain gap.
6.3. Texturing
Since the base features are obtained from image models (see Sec. 3) and our pointwise encoder can process points sampled from a mesh as easily as pixels sampled from an image, we can establish correspondences between a 2D image and a 3D mesh. We demonstrate this by texturing 3D meshes from a masked target image and individually assign each vertex a color from the image pixel that maximizes the mutual feature similarity (Eq. 1) (see Fig. 13 and Appendix D.3). We observe that our features produce a more coherent mapping leading to a better preservation of the source appearance when compared to the Diff3F base features. In Fig. 14, we further show that textures can also be effectively transferred between two 3D shapes using a combination our surface-aware features with Functional Maps like in Sec. 5.4.
Figure 13. Texturing of 3D meshes from SMPLeval and TOSCA, based on a 2D image generated with ChatGPT. The appearance is transferred by establishing correspondence between the image features and 3D mesh features. In contrast to Diff3F, our surface-aware features represent the input image more faithfully.
Figure 14. Texturing of 3D meshes based on a source 3D mesh. The appearance is transferred based on correspondences established by combining our surface-aware features with Functional Maps. The data originate from SMPLeval, SMPLitex (Casas and Comino-Trinidad, 2023), and DeformingThings4D (Li et al., 2021).
7. Discussion
Limitations and Future Work
Our method inherits limitations connected to the extraction of the base features. Specifically, the extraction of Diff3F features (Dutt et al., 2024) takes several minutes per mesh and its vision model is sensitive to rendering artifacts or upside-down mesh orientations. We expect that advances in rendering of point representations increase the applicability across representations (Kerbl et al., 2023). Furthermore, our method cannot establish a consistent partitioning for objects that are both geometrically and semantically isotropic (e.g., a round table). Hence, while our embedding separates human legs following the body’s notion of front and rear, it cannot do so for table legs. However, this is not an issue for applications such as shape morphing (Sun et al., 2024). Lastly, our method relies on consistency of geodesic distances between semantically distinct parts, and therefore it will be affected by geodesic shortcuts for partially blended parts in noisy 3D reconstructions (e.g., touching hands).
Beyond 3D alignment, our methodology could inspire 3D-to-2D pose estimation (Kanazawa et al., 2018; Peng et al., 2019), articulated 3D reconstruction (Yao et al., 2022; Uzolas et al., 2023), automated rigging (Xu et al., 2020) or 2D-to-3D uplifting (Poole et al., 2022; Liu et al., 2023), where our features could support more view-consistent representations. Finally, an interesting topic for future research is the development of foundational features using massive datasets, such as Objaverse (Deitke et al., 2023).
Conclusion
We have introduced novel surface-aware features for 3D shape matching that disambiguate intra-class instances among semantic features derived from pre-trained 2D vision models. Our descriptors have proven effective in distinguishing instances of the same semantic class and they generalize even when trained on a limited number of 3D shapes. Furthermore, our contrastive loss facilitates easy integration in future unsupervised methods which reduces data labeling effort. Consequently, our method is a promising building block toward adapting pre-trained 2D models to 3D tasks.
Acknowledgements.
This work was partially supported by the Convergence AI Immersive Tech Lab at TU Delft.
Appendix A Additional implementation details
Here, we provide additional details to reproduce our experiments.
A.1. Diff3F
We use the authors’ code to compute the Diff3F features (Dutt et al., 2024) for our base features and as a baseline method for our comparisons.
Since we observed that the camera poses used for sampling in the TOSCA dataset are biased towards a specific up-direction, we modified the code to flip the coordinate system for only this dataset. Doing so yields $\approx 10$% increase in correspondence accuracy in TOSCA for both our method and the Diff3F baseline, when compared to the numbers reported in the Diff3F paper (Dutt et al., 2024). An alternative solution could be a modification of the camera sampling algorithm itself.
A.2. Ablations
Figure 15. Ablation on the anchor count AA on SHREC’19. In terms of the correspondence error, our method improves upon Diff3F already starting from $A=1$.
Number of Anchors
We train our method for different anchor counts $A$ with a constant two-hour training budget per model. In Fig. 15, we observe a low sensitivity to the anchor count. Due to repeated random sampling over the course of training, even $A=1$ outperforms the correspondence error of Diff3F. Ultimately, we opt for $A=100$ in all our experiments, as it balances computation cost and matching performance well. Note that neither the parameter $A$ nor any computation of geodesic distances in generally are used during inference.
Losses
Utilizing the geodesic distance as a supervision signal for embeddings has been explored in related works  (Jiang et al., 2023; Chen et al., 2025). However, in our ablations, we show that we achieve superior results with our formulation. In this section, we discuss the key differences.
Relative Geodesic Loss (RGL)
Based on two points, $v_{p}$ and $v_{q}$, the Relative Geodesic Loss (Jiang et al., 2023) minimizes the difference between the geodesic distance $d^{S}$ and the Euclidean embedding distance $d^{E}$ of those two points, relative to the surface distance:
(4) $L_{RGL}=\sum_{i}\sum_{(p,q)\in S_{i}}\frac{|d_{i}^{E}(v_{p},v_{q})-d^{S}(v_{p},v_{q})|^{2}}{d^{S}(v_{p},v_{q})^{2}}.$
The normalization term is introduced to prioritize local distance preservation. We do not utilize this normalization term, because the base features struggle to disambiguate samples that are far away on the surface but close in feature space, such as “left hand” and “right hand”.
Naive Geodesic Loss (NGL)
NGL is discussed by Jiang et al. (2023) but not used for training, as the authors state that it might hamper local distance preservation. Indeed, in our ablations, it achieves worse results in terms of correspondence accuracy when compared to RGL. It is identical to RGL but it omits the normalization term:
(5) $L_{NGL}=\sum_{i}\sum_{(p,q)\in S_{i}\in[n_{i}]^{2}}|d_{i}^{E}(v_{p},v_{q})-d^{S}(v_{p},v_{q})|^{2}$
While not actually utilized in their work, the NGL formulation is the closest of the three to our own formulation. However, our choice of a hyperspherical rather than Euclidean embedding space in combination with our autoencoder setup achieves notably better results.
Geometric Similarity Loss (GSL)
A concurrent work proposes to maximize the local angular similarity between geodesic distances and Euclidean feature vectors for a set of neighbors for each point (Chen et al., 2025). Crucially, the similarity is only maximized for a set of $k$ point pairs nearest in the embedding space. The loss is a cosine metric between a vector of geodesic distances $\mathbf{m}_{i}\in\mathbb{R}^{k}$ and a vector of Euclidean embedding distances $\mathbf{d}_{i}\in\mathbb{R}^{k}$:
(6) $\mathcal{L}_{\text{GSL}}=\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{\mathbf{d}_{i}\cdot\mathbf{m}_{i}}{\|\mathbf{d}_{i}\|\|\mathbf{m}_{i}\|}\right).$
This restricts GSL supervision to a fixed neighborhood size and potentially limits disambiguation of features that are close in feature space but not among the $k$ nearest neighbors. In contrast, our method follows a global approach by sampling anchors based on a furthest point sampling.
Conclusion
Our method differs from recent and concurrent works utilizing vision-based features for 3D shape matching in three main aspects: 1) we follow a global approach when enforcing distances in the embedding space; 2) our embedding space is hyperspherical and it only encodes angular information; 3) in the context of the whole pipeline, we solely rely on intrinsic properties.
Appendix B One Shot Pose Transfer
We train an MLP to model the deformation between the paired input source mesh $\mathcal{M}_{train}^{src}$ and the output target mesh $\mathcal{M}_{train}^{tgt}$ obtained from SMPLevaland thus not used for training of our features. The paired training meshes share the same identity $\beta$ but they differ in poses $\theta$ such that
$\mathcal{M}_{train}^{src}$ $:=SMPL(\beta_{src},\theta_{src})$
$\mathcal{M}_{train}^{tgt}$ $:=SMPL(\beta_{src},\theta_{tgt}).$
We train an MLP $\mathcal{M}_{train}^{tgt}=\mathcal{J}(\mathcal{M}_{train}^{src},\mathcal{S_{M}^{\text{{src}}}})$ to produce the target pose mesh $\mathcal{M}_{train}^{tgt}$ conditioned on the source mesh $\mathcal{M}_{train}^{src}$ and its surface-aware features $\mathcal{S_{M}^{\text{{src}}}}$. These features are produced by our pre-trained general encoder $\mathcal{E}(.)$ from Sec. 5 without any further fine-tuning. Alternatively, other features are used for comparisons.
During training, $\mathcal{J}(.)$ learns a per-face residual function  (Muralikrishnan et al., 2024) to model a Neural Jacobian Field (Aigerman et al., 2022) while being supervised by MSE between the predicted re-posed mesh and the target $\mathcal{M}_{train}^{tgt}$.
Crucially, $\mathcal{J}(.)$ is defined on a per-face basis, which means that the input meshes used in test time do not need to have the same connectivity as the training mesh pair. In quantitative comparisons, we apply the learned mapping $\mathcal{J}(.)$ to input meshes of unseen identities $\mathcal{M}_{test}^{src}:=SMPL(\beta_{test},\theta_{src})$ and compare against their ground truths $\mathcal{M}_{test}^{tgt}:=SMPL(\beta_{test},\theta_{tgt})$. For other shape classes, we provide qualitative comparisons.
Appendix C Pose Alignment
We establish correspondences between two input shapes based on the feature cosine similarity $\phi(.)$ (Sec. 3), such that each point $\mathbf{x}_{i}^{\mathcal{S}}$ in the source shape is assigned a target point $\mathbf{x}_{i}^{\mathcal{T}}$ in the target shape. Next, we align the source to the target by minimizing the following L1 loss:
[…]
For the first half of the optimization steps, we only optimize the root rotation $\mathbf{R}$, the root translation $\mathbf{t}$, and the scale $s$, which roughly rigidly aligns the meshes. In the second half, we additionally optimize the rotation $\mathbf{R}_{b}$ of each kinematic bone $b$. The parameters are optimized based on a gradient-descent for $4000$ iterations, which takes approximately 30 seconds for a static pose.
Furthermore, we found it beneficial to use an as-rigid-as-possible regularization (Sorkine and Alexa, 2007), which penalizes the deviation between the initial edge lengths of the mesh $\delta^{init}_{e}$ and the current edge length $\delta_{e}$ for each edge $e$:
(8) $\mathcal{L}_{arap}=\frac{1}{E}\sum_{e}^{E}|\delta_{e}^{init}-\delta_{e}|.$
When fitting an animation as a pose sequence, we optimize the pose parameters for each time step $t$. Furthermore, we apply pointwise temporal smoothing for neighboring frames:
[…]
The final loss is $\mathcal{L}_{pose}=w_{p}\mathcal{L}_{point}+w_{a}\mathcal{L}_{arap}+w_{s}\mathcal{L}_{smooth}$ with $w_{p}=w_{a}=w_{s}=1$ for animations and $w_{s}=0$ otherwise.
Appendix D Additional results
D.1. Qualitative Results on SHREC’20
Fig. 16 presents additional results for the SHREC’20 dataset. As the dataset only provides $\approx 50$ correspondences for each shape pair, we display the predicted correspondences without dense ground-truth labels. However, we find that our features generally produce smoother correspondences (e.g., bottom left) and a better separation of legs (e.g., the second to last row on the right).
Figure 16. Qualitative results on SHREC’20. Our features result on average in a more accurate smooth mapping between different limbs.
D.2. Qualitative Results on ShapeNet
In Fig. 17, we show additional results for chairs and airplanes from ShapeNet (Chang et al., 2015). Since no dense ground truth labels are available, we show the predicted correspondences alone. We find that our surface-aware features achieve results better than the Diff3F baseline when separating the chair legs (see the top left row) and the airplane wings (see the top right row).
Figure 17. Qualitative results on ShapeNet (Chang et al., 2015). Our features result on average in a more accurate smooth mapping between chair legs and wings.
D.3. Texturing
We provide additional examples of 2D-to-3D texturing based on our own features in Fig. 18.
Figure 18. Texturing of 3D meshes from SMPLeval and TOSCA, based on a 2D image generated with ChatGPT. The appearance is transferred by establishing correspondence between the image features and 3D mesh features. Our method performs well even on incomplete meshes (notice the bear in the second row).
Table 3. Correspondence metrics measured for human and animal test shapes. The Specialized encoder was trained on a join set of humans and animals following our point-to-point correspondence experiment procedure in Sec. 5.4, and thus, the values match Tbl. 1 in the paper (see Ours). All-shape encoder was trained on a larger more generalized training set covering humans and animals but also chairs and airplanes.
Encoder SHREC’19 TOSCA SHREC’20
Specialized err ↓ 0.43±0.76 1.65±2.15 23.89±8.90
acc ↑ 28.78±9.30 29.35±14.53 73.97±26.47
All-shape err ↓ 0.56±1.03 1.62±2.08 4.37±9.47
acc ↑ 27.53±9.47 27.45±14.56 70.33±24.86
D.4. Using a Shared Encoder for All Shapes
In Sec. 5, we train a shared encoder for (SURREAL) and animals (SMAL) shapes as well as separate encoders for chair and airplanes (ShapeNet). Here, we follow the same procedure and train a new single shared encoder on a union of all these shapes and test it on human and animal shapes as in the paper Tbl. 1 to assess further generalization of our approach. The correspondence metrics in Tbl. 3 show that the all-shape encoder generally slightly under-performs the specialized encoders but it still improves upon the baselines (see Tbl. 1 in the paper). TOSCA is an exception, as the larger combined train set marginally reduces the error.
D.5. Ablating the Number of Training Shapes
In Tbl. 4, we explore how the number of training shapes affects the correspondence error with the same fixed validation set and training policy as in the main experiments. We vary the number of training samples while retaining a constant animal-to-human shape ratio. We find that just 2 training samples already decrease the error when compared to Diff3F. As expected, additional samples lead to further improvements for SHREC’19 and TOSCA. This trend is more subtle for SHREC’20.
Table 4. Training set size (columns) vs. correspondence error.
Diff3f 2 10 50 94
SHREC’19 1.69±1.44 1.32±1.22 1.31±1.31 0.48±0.85 0.43±0.76
TOSCA 4.51±5.48 3.75±3.50 2.60±2.74 1.84±2.47 1.65±2.15
SHREC’20 5.34±10.22 3.89±8.49 4.05±9.90 3.96±9.33 3.89±8.90